OpenStudy (anonymous):

Solve. 4x = –36 –9 –4 4 9

OpenStudy (anonymous):

-9

OpenStudy (anonymous):

4x=-36 x=-36/4 x=-9

OpenStudy (arabpride):

4x = -36 <divide from both sides> 4x = -36 /4 /4 x = -9

OpenStudy (anonymous):

thanks so much :)

OpenStudy (xxdragoonxx):

-36 divide by 4 = -9